OpenStudy (anonymous):

Find the roots of the equation x^2 = 30 -13x

OpenStudy (nathan917):

Simplifying x2 = 30 + -13x Solving x2 = 30 + -13x Solving for variable 'x'. Reorder the terms: -30 + 13x + x2 = 30 + -13x + -30 + 13x Reorder the terms: -30 + 13x + x2 = 30 + -30 + -13x + 13x Combine like terms: 30 + -30 = 0 -30 + 13x + x2 = 0 + -13x + 13x -30 + 13x + x2 = -13x + 13x Combine like terms: -13x + 13x = 0 -30 + 13x + x2 = 0 Factor a trinomial. (-15 + -1x)(2 + -1x) = 0 Subproblem 1 Set the factor '(-15 + -1x)' equal to zero and attempt to solve: Simplifying -15 + -1x = 0 Solving -15 + -1x = 0 Move all terms containing x to the left, all other terms to the right. Add '15' to each side of the equation. -15 + 15 + -1x = 0 + 15 Combine like terms: -15 + 15 = 0 0 + -1x = 0 + 15 -1x = 0 + 15 Combine like terms: 0 + 15 = 15 -1x = 15 Divide each side by '-1'. x = -15 Simplifying x = -15 Subproblem 2 Set the factor '(2 + -1x)' equal to zero and attempt to solve: Simplifying 2 + -1x = 0 Solving 2 + -1x = 0 Move all terms containing x to the left, all other terms to the right. Add '-2' to each side of the equation. 2 + -2 + -1x = 0 + -2 Combine like terms: 2 + -2 = 0 0 + -1x = 0 + -2 -1x = 0 + -2 Combine like terms: 0 + -2 = -2 -1x = -2 Divide each side by '-1'. x = 2 Simplifying x = 2 Solution x = {-15, 2}

OpenStudy (nathan917):

Pls fan if you think you will have anymore.

OpenStudy (anonymous):

thank you so much!